OpenStudy (anonymous):

A food store makes a 11–pound mixture of peanuts, almonds, and raisins. The cost of peanuts is $1.50 per pound, almonds cost $3.00 per pound, and raisins cost $1.50 per pound. The mixture calls for twice as many peanuts as almonds. The total cost of the mixture is $21.00. How much of each ingredient did the store use? 3 lbs peanuts, 6 lbs almonds, 2 lbs raisins 8 lbs peanuts, 1 lb almonds, 2 lbs raisins 6 lbs peanuts, 3 lbs almonds, 2 lbs raisins 8 lbs peanuts, 2 lbs almonds, 1 lbs raisins

jimthompson5910 (jim_thompson5910):

Start off by defining variables

jimthompson5910 (jim_thompson5910):

let x = amount of peanuts (in pounds), y = amount of almonds (in pounds) and z = amount of raisins (in pounds)

OpenStudy (anonymous):

okey (:

jimthompson5910 (jim_thompson5910):

If "A food store makes a 11–pound mixture of peanuts, almonds, and raisins" then what does this mean?

OpenStudy (anonymous):

That they add up to 11 pounds haha

jimthompson5910 (jim_thompson5910):

good, so x y and z add to 11, which means x+y+z = 11

jimthompson5910 (jim_thompson5910):

that's one equation

OpenStudy (anonymous):

so x+y+z=11 :o

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

But then we have to factor in how much they cost..

jimthompson5910 (jim_thompson5910):

We also know that "The cost of peanuts is $1.50 per pound, almonds cost $3.00 per pound, and raisins cost $1.50 per pound" So x lbs of peanuts cost 1.5x dollars, y lbs of almonds cost 3y dollars, and z lbs of raisins cost 1.5z dollars They total 1.5x + 3y + 1.5z dollars

jimthompson5910 (jim_thompson5910):

this total cost is given to be $21 so 1.5x + 3y + 1.5z = 21

OpenStudy (anonymous):

so like, 1.50x+3.00y+1.50=21..

OpenStudy (anonymous):

Haha Okey you already said that x.x my bad

OpenStudy (anonymous):

1.50z* I meant btw haha. I get that indeed.

jimthompson5910 (jim_thompson5910):

close, you mean to say 1.50x+3.00y+1.50z=21 but you got it

OpenStudy (anonymous):

And don't forget twice the peanuts!

jimthompson5910 (jim_thompson5910):

that's your 2nd equation

jimthompson5910 (jim_thompson5910):

twice the peanuts as what?

OpenStudy (anonymous):

it says, the recipe calls for twice as many peanuts as almonds O.o

jimthompson5910 (jim_thompson5910):

good, so the equation for that is?

OpenStudy (anonymous):

ermm.. I dunno lol.. x(times2y)+y+z=11 O.O I have no idea.

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (anonymous):

knew it was wrong x.x

jimthompson5910 (jim_thompson5910):

Amount of peanuts = 2 times (amount of almonds)

jimthompson5910 (jim_thompson5910):

y = 2 times x or just y = 2x

OpenStudy (anonymous):

oh pshh I should have thought of that x.x

jimthompson5910 (jim_thompson5910):

so you have these 3 equations x+y+z = 11 1.5x + 3y + 1.5z = 21 y = 2x

jimthompson5910 (jim_thompson5910):

you can use them to solve for x, y and z

jimthompson5910 (jim_thompson5910):

I'll get you started

OpenStudy (anonymous):

Lol I can't because I'm stupid, But smart people definitely can haha

jimthompson5910 (jim_thompson5910):

x+y+z = 11 x+2x+z = 11 3x+z = 11 --------------------- 1.5x + 3y + 1.5z = 21 1.5x + 3(2x) + 1.5z = 21 1.5x + 6x + 1.5z = 21 7.5x + 1.5z = 21

jimthompson5910 (jim_thompson5910):

so after plugging in y = 2x, you now have the new system 3x+z = 11 7.5x + 1.5z = 21

OpenStudy (anonymous):

O.o That all made absolutely no sense to me but it seems to for you and that's good enough for me

jimthompson5910 (jim_thompson5910):

where are you lost?

jimthompson5910 (jim_thompson5910):

I took x+y+z = 11 plugged in y = 2x (ie replaced every 'y' with '2x') to get x+2x+z = 11 then I combined like terms to get 3x+z = 11

jimthompson5910 (jim_thompson5910):

I did the same with 1.5x + 3y + 1.5z = 21 to get 7.5x + 1.5z = 21

OpenStudy (anonymous):

oh okay ! I see now, sorry it's really late and I'm super tired :/

jimthompson5910 (jim_thompson5910):

so you have 3x+z = 11 7.5x + 1.5z = 21 now

jimthompson5910 (jim_thompson5910):

solve the first equation for z to get this 3x+z = 11 3x+z - 3x = 11 - 3x z = 11 - 3x

OpenStudy (anonymous):

Alrighty

jimthompson5910 (jim_thompson5910):

Since z = 11 - 3x, you can plug it into the other equation and solve for x 7.5x + 1.5z = 21 7.5x + 1.5(11 - 3x) = 21 Solve for x

jimthompson5910 (jim_thompson5910):

oh wait, i made a variable mistake...i switched x and y one sec

jimthompson5910 (jim_thompson5910):

Let me fix real quick

OpenStudy (anonymous):

I was thinking that haha x) ..

OpenStudy (anonymous):

uhh ? your comment disappeared .. :o

jimthompson5910 (jim_thompson5910):

it should be x = 2y, not y = 2x so x+y+z = 11 2y+y+z = 11 3y + z = 11 z = 11 - 3y --------------------------- 1.5x + 3y + 1.5z = 21 1.5(2y) + 3y + 1.5z = 21 3y + 3y + 1.5z = 21 6y + 1.5z = 21 =========================== 6y + 1.5z = 21 6y + 1.5(11 - 3y) = 21 Solve that last equation for y

jimthompson5910 (jim_thompson5910):

sry

jimthompson5910 (jim_thompson5910):

Solve 6y + 1.5(11 - 3y) = 21 for y

OpenStudy (anonymous):

haha it's okay

OpenStudy (anonymous):

Aw man I have to go to bed soon I'm dead tired x.x

jimthompson5910 (jim_thompson5910):

What do you get when you solve that equation for y

OpenStudy (anonymous):

ergh I dunno x.x I'm brain dead

jimthompson5910 (jim_thompson5910):

Then give it a break, get some sleep and you'll tackle it in the morning

OpenStudy (anonymous):

Oh no I keep losing connection >.<